# Context-aware Inductive Knowledge Graph Completion with Latent Type Constraints and Subgraph Reasoning

### v2

@@ Introduction
Graph Neural Networks (GNNs) (Song, Zhang, and King 2023b, a) are popular in natural language processing for modeling relationships between entities (Ma et al. 2023). GNN-based methods such as CompGCN (Vashishth et al. 2020), RGCN (Schlichtkrull et al. 2018), and WGCN (Zhao et al. 2021) embed entities in a KG by iteratively aggregating features from the local neighborhood. Nevertheless, these approaches struggle to produce meaningful embeddings for newly emerged entities with few links to existing ones, and are not applicable to entirely new graphs (Zha, Chen, and Yan 2022). To conform to the inductive setting, GraIL (Teru, Denis, and Hamilton 2020), and TACT (Chen et al. 2021) embed entities in the local subgraph with their distances to the head and tail entities of the query triple. However, such an embedding approach fails to distinguish different entities that share the same relative position. Consequently, it cannot perform well when the subgraph of query triple is large.
@@
$\mathcal{T}_{h_{q}}=\operatorname*{arg\,max}_{(h_{q},r,t)\in\mathcal{T}_{\text{train}}}cos(f_{\text{bge}}(h_{q},r,t),f_{\text{bge}}(h_{q},r_{q},t_{q})),$ (4)
$\mathcal{T}_{t_{q}}=\operatorname*{arg\,max}_{(h,r,t_{q})\in\mathcal{T}_{\text{train}}}cos(f_{\text{bge}}(h,r,t_{q}),f_{\text{bge}}(h_{q},r_{q},t_{q})),$ (5)
@@
$\vspace{-0.25cm}\mathcal{L}_{\text{SR}}=$ $-\sum_{(h,r,t)\in\mathcal{T}_{\text{train}}}\log{p(\hat{y}=\text{`Y'}|\mathcal{P}_{(h_{q},t_{q})},\mathcal{T}_{h_{q}},\mathcal{T}_{t_{q}};\Theta)}$
$-\sum_{(h,r,t)\notin\mathcal{T}_{\text{train}}}\log{p(\hat{y}=\text{`N'}|\mathcal{P}_{(h_{q},t_{q})},\mathcal{T}_{h_{q}},\mathcal{T}_{t_{q}};\Theta)}.\vspace{-0.25cm}$ (6)
@@
Transductive Inductive
Metric Method WN18RR FB15k-237 NELL-995 WN18RR FB15k-237 NELL-995
MRR RuleN 0.669 0.674 0.736 0.780 0.462 0.710
GRAIL 0.676 0.597 0.727 0.799 0.469 0.675
MINERVA 0.656 0.572 0.592 - - -
TuckER 0.646 0.682 0.800 - - -
KG-BERT - - - 0.547 0.500 0.419
BERTRL 0.683 0.695 0.781 0.792 0.605 0.808
KRST 0.899 0.720 0.800 0.890 0.716 0.769
APST 0.902 0.774 0.801 0.908 0.764 0.769
CATS (TAR) 0.956 0.812 0.836 0.937 0.834 0.750
CATS (SR) 0.972 0.829 0.869 0.992 0.875 0.906
CATS (full) 0.978 0.843 0.885 0.982 0.882 0.861
Hits@1 RuleN 0.646 0.603 0.636 0.745 0.415 0.638
GRAIL 0.644 0.494 0.615 0.769 0.390 0.554
MINERVA 0.632 0.534 0.553 - - -
TuckER 0.600 0.615 0.729 - - -
KG-BERT - - - 0.436 0.341 0.244
BERTRL 0.655 0.620 0.686 0.755 0.541 0.715
KRST 0.835 0.639 0.694 0.809 0.600 0.649
APST 0.839 0.694 0.698 0.837 0.643 0.663
CATS (TAR) 0.922 0.726 0.745 0.888 0.744 0.624
CATS (SR) 0.951 0.752 0.792 0.984 0.804 0.849
CATS (full) 0.962 0.776 0.820 0.965 0.805 0.783
Experiments
Datasets and Evaluation Metrics
We evaluate our proposed method on three widely adopted benchmark KGs WN18RR, FB15k-237, and NELL-995 with their transductive and inductive subsets. Following the convention of (Zha, Chen, and Yan 2022; Su et al. 2023, 2024), we evaluate each query triple with $1$ ground truth entity and $49$ negative candidate entities. For a fair comparison, we use the same dataset splits and negative triples provided by (Zha, Chen, and Yan 2022). The detailed statistics of datasets are available in the technical appendix. We score the plausibility of the query triple with each candidate entity and rank them in descending order. The performance of our model is evaluated based on two metrics: Mean Reciprocal Rank (MRR) and Hits@$1$.
Baselines and Experiment Settings
[…]
Main Results
We evaluate the proposed CATS framework on the three datasets in transductive and inductive settings. The “TAR” and “SR” variants of CATS severally utilize the probabilities generated by corresponding modules to score and rank each candidate entity. Experimental Results in Table 7 demonstrate that the CATS (full) significantly and consistently outperforms all baseline methods. Most notably, CATS (full) achieves double-digit absolute Hits@1 improvements of $12.8\%$, $16.2\%$, and $12.0\%$ on the WN18RR, FB15k-237, and NELL-995 datasets under an inductive setting. Correspondingly, the improvements in transductive scenarios, namely $12.3\%$, $8.2\%$, and $12.2\%$, are also remarkable.
Transductive Inductive
WN18RR FB15k-237 NELL-995 WN18RR FB15k-237 NELL-995
MINERVA 0.125 0.268 0.198 0.364 0.182 0.322 - - - - - -
TuckER 0.258 0.448 0.457 0.601 0.436 0.577 - - - - - -
KG-BERT - - - - - - 0.471 0.525 0.431 0.460 0.406 0.406
Table 2: Few-shot KGC results on WN18RR, FB15k-237 and NELL-995
Ablation Studies
Table 8 shows the experiment results on the three datasets with $1000$ and $2000$ triples in the training graph. In general, CATS achieves significant improvements for $10$ out of $12$ cases in terms of MRR compared to state-of-the-art methods.  [^4] Results in terms of Hits@$1$ are reported in the appendix. For the remaining cases, the performance gap to the best baseline method is marginal. Considering the three variants of CATS, TAR outperforms SR in transductive settings, while SR performs better in inductive scenarios. The disparity in performance can be attributed to the following reasons: Reducing the number of triples in the training graph significantly decreases the average degree of each entity. Consequently, it becomes challenging to sample a sufficient number of neighboring facts and reasoning paths for entities in the query triple, which diminishes the effectiveness of the SR variant. However, in most KGs, the number of relations is considerably smaller than the number of entities ($|\mathcal{R}|<<|\mathcal{E}|$). Hence, we are still able to retrieve sufficient triples with the same relation from the training graph, which sustains the desirable performance of the TAR variant, and ensures the robustness of the the proposed framework. In inductive scenarios, neighboring facts and reasoning paths are sampled from the supplementary test graph. Hence, reducing training triples has subtle negative impacts on the effectiveness of the SR variant.
One may notice that CATS does not achieve stat-of-the-art performance on the NELL-995 dataset in transductive settings. This is attributed to the higher number of triples in the NELL-995 training graph. Selecting the subset of $1000$ or $2000$ triples results in an extremely sparse graph structure. Without external knowledge, the training graph may fail to provide sufficient support for the inference of certain triples, thereby lowering the performance ceiling. Moreover, KRST (Su et al. 2023) and APST (Su et al. 2024) take advantage of extra knowledge from entity descriptions, allowing them to catch up with CATS.  [^5] We adopt entity and relation labels as the sole textual context.
Transductive Inductive
Configuration WN FB NE WN FB NE
CATS (SR) 0.972 0.829 0.869 0.992 0.875 0.906
- w/ NF only 0.962 0.824 0.861 0.971 0.873 0.895
- w/ RP (filt.) only 0.960 0.800 0.773 0.968 0.851 0.829
- w/ RP only 0.954 0.801 0.769 0.965 0.844 0.821
Transductive Inductive
LLM & Config. WN FB NE WN FB NE
Previous SOTA 0.902 0.774 0.801 0.908 0.764 0.808
Llama-3-8B (full) 0.965 0.802 0.867 0.956 0.862 0.837
Qwen2-1.5B (full) 0.966 0.804 0.877 0.948 0.814 0.835
Qwen2-7B (comb.) 0.967 0.830 0.861 0.985 0.859 0.885
Qwen2-7B (full) 0.978 0.843 0.885 0.982 0.882 0.861
- w/o. TAR & SR 0.947 0.800 0.811 0.901 0.814 0.710
- w/o. SFT 0.225 0.197 0.177 0.207 0.179 0.208
- w/o. all 0.199 0.140 0.156 0.174 0.144 0.138
Table 4: Transductive and Inductive KGC performance (in MRR) with different LLMs and configurations. Here Qwen2-7B (full) refers to CATS (full).
Conclusion
Appendix A Technical Appendix
@@
Dataset Train Splits $|Y|$ $|N|$ Average Prompt Length
FB15k-237 full 8490 50940 1034.36
1000 2054 12324 777.99
2000 4016 24096 870.82
NELL-995 full 16438 98628 1082.78
1000 2040 18360 778.29
2000 4022 36198 884.75
WN18RR full 10820 64920 2038.06
2000 4004 24024 1656.73
1000 2002 12012 1544.91
Table 6: Statistics of supervised fine-tuning instructions across different datasets and training splits, including the number of positive and negative samples and the average length of the prompts.
@@
Table 7: Transductive and inductive KGC results on WN18RR, FB15k-237 and NELL-995 datasets.
@@
Table 8: Few-shot KGC results on WN18RR, FB15k-237 and NELL-995 datasets
@@
Table 10: Transductive and Inductive KGC performance (in MRR) with different LLMs and configurations. Here, Qwen2-7B (full) refers to CATS (full).
LLM Config. WN18RR FB15k-237 NELL-995 Average
Qwen2-7B TAR 0.830 0.395 0.407 0.544
SR 1.091 0.714 0.876 0.894
Total 1.921 1.109 1.282 1.437
LLama3-8B TAR 0.862 0.412 0.414 0.563
SR 1.165 0.718 0.901 0.928
Total 2.026 1.130 1.315 1.490
Qwen2-1.5B TAR 0.292 0.130 0.144 0.189
SR 0.438 0.240 0.279 0.319
Total 0.730 0.369 0.423 0.507
Table 11: Time cost per ranking query (in seconds) for different LLMs and configurations across WN18RR, FB15k-237, and NELL-995 datasets on 2 GPUs. The average values are calculated for each LLM and overall across datasets.
WN18RR RB15k-237 NELL-995
Qwen2-7B 2h12min 1h35min 3h19min
Table 12: Time cost of supervised fine-tuning on WN18RR, RB15k-237, and NELL-995 datasets. Note: Results may vary in different experimental configurations.

### v3

@@ Introduction
Graph neural networks (GNNs) (Song, Zhang, and King 2023a, b) are popular in natural language processing for modeling relationships between entities (Ma et al. 2023). GNN-based methods such as CompGCN (Vashishth et al. 2020), RGCN (Schlichtkrull et al. 2018), and WGCN (Zhao et al. 2021) embed entities in a KG by iteratively aggregating features from the local neighborhood. Nevertheless, these approaches struggle to produce meaningful embeddings for newly emerged entities with few links to existing ones, and are not applicable to entirely new graphs (Zha, Chen, and Yan 2022). To conform to the inductive setting, GraIL (Teru, Denis, and Hamilton 2020), and TACT (Chen et al. 2021) embed entities in the local subgraph with their distances to the head and tail entities of the query triple. However, such an embedding approach fails to distinguish different entities that share the same relative position. Consequently, it cannot perform well when the subgraph of query triple is large. RED-GNN (Zhang and Yao 2022) and Adaprop (Zhang et al. 2023) enhance the message-passing mechanisms through progressive and adaptive propagation. Still, these methods fail to address the suboptimal performance of GNNs on sparse graph structures.
@@
$\mathcal{T}_{h_{q}}=\operatorname*{arg\,max}_{(h_{q},r,t)\in\mathcal{T}_{\text{train}}}\cos(f_{\text{bge}}(h_{q},r,t),f_{\text{bge}}(h_{q},r_{q},t_{q})),$ (4)
$\mathcal{T}_{t_{q}}=\operatorname*{arg\,max}_{(h,r,t_{q})\in\mathcal{T}_{\text{train}}}\cos(f_{\text{bge}}(h,r,t_{q}),f_{\text{bge}}(h_{q},r_{q},t_{q})),$ (5)
@@
$\mathcal{L}_{\text{SR}}=$ $-\sum_{(h,r,t)\in\mathcal{T}_{\text{train}}}\log{p(\hat{y}=\text{`Y'}|\mathcal{P}_{(h_{q},t_{q})},\mathcal{T}_{h_{q}},\mathcal{T}_{t_{q}};\Theta)}$
$-\sum_{(h,r,t)\notin\mathcal{T}_{\text{train}}}\log{p(\hat{y}=\text{`N'}|\mathcal{P}_{(h_{q},t_{q})},\mathcal{T}_{h_{q}},\mathcal{T}_{t_{q}};\Theta)}.$ (6)
@@
Table 3: Few-shot KGC results on WN18RR, FB15k-237 and NELL-995
Experiments
Datasets and Evaluation Metrics
We evaluate our proposed method on three widely adopted benchmark KGs WN18RR (Miller 1995), FB15k-237 (Bollacker et al. 2008), and NELL-995 (Carlson et al. 2010) with their transductive and inductive subsets. Following the convention of (Zha, Chen, and Yan 2022; Su et al. 2023, 2024), we evaluate each query triple with $1$ ground truth entity and $49$ negative candidate entities. For a fair comparison, we use the same dataset splits and negative triples provided by (Zha, Chen, and Yan 2022). The detailed statistics of datasets are available in Table 1. We score the plausibility of the query triple with each candidate entity and rank them in descending order. The performance of our model is evaluated based on two metrics: Mean Reciprocal Rank (MRR) and Hits@$1$.
Baselines and Experiment Settings
[…]
Main Results
We evaluate the proposed CATS framework on the three datasets in transductive and inductive settings. The “TAR” and “SR” variants of CATS severally utilize the probabilities generated by corresponding modules to score and rank each candidate entity. Experimental Results in Table 2 demonstrate that the CATS (full) significantly and consistently outperforms all baseline methods. Most notably, CATS (full) achieves absolute Hits@1 improvements of $12.8\%$, $16.2\%$, and $6.8\%$ on the WN18RR, FB15k-237, and NELL-995 datasets under an inductive setting. Correspondingly, the improvements in transductive scenarios, namely $12.3\%$, $8.2\%$, and $9.1\%$, are also remarkable.
Ablation Studies
Table 3 shows the experiment results on the three datasets with $1000$ and $2000$ triples in the training graph. In general, CATS achieves significant improvements for $10$ out of $12$ cases in terms of MRR compared to state-of-the-art methods. For the remaining cases, the performance gap to the best baseline method is marginal. Considering the three variants of CATS, TAR outperforms SR in transductive settings, while SR performs better in inductive scenarios. The disparity in performance can be attributed to the following reasons: Reducing the number of triples in the training graph significantly decreases the average degree of each entity. Consequently, it becomes challenging to sample a sufficient number of neighboring facts and reasoning paths for entities in the query triple, which diminishes the effectiveness of the SR variant. However, in most KGs, the number of relations is considerably smaller than the number of entities ($|\mathcal{R}|<<|\mathcal{E}|$). Hence, we are still able to retrieve sufficient triples with the same relation from the training graph, which sustains the desirable performance of the TAR variant, and ensures the robustness of the the proposed framework. In inductive scenarios, neighboring facts and reasoning paths are sampled from the supplementary test graph. Hence, reducing training triples has subtle negative impacts on the effectiveness of the SR variant.
One may notice that CATS does not achieve stat-of-the-art performance on the NELL-995 dataset in transductive settings. This is attributed to the higher number of triples in the NELL-995 training graph. Selecting the subset of $1000$ or $2000$ triples results in an extremely sparse graph structure. Without external knowledge, the training graph may fail to provide sufficient support for the inference of certain triples, thereby lowering the performance ceiling. For the same reason, GNN-based methods such as RED-GNN (Zhang and Yao 2022) and Adaprop (Zhang et al. 2023) exhibit significant performance drops in few-shot scenarios. Moreover, KRST (Su et al. 2023) and APST (Su et al. 2024) take advantage of extra knowledge from entity descriptions, allowing them to catch up with CATS.
@@
Table 5: Transductive and Inductive KGC performance (in MRR) with different LLMs and configurations.
Conclusion
Acknowledgment
The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 14222922, RGC GRF 2151185).